A Complete Bibliography of Euleriana

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

(maa + nbb) = cd(mcc + ndd) [Ehl23]. π [CW23].

Ad [Hea23]. Aequationem [Hea22]. analyse [GS21]. Analytical [Huf23]. Answer [Ayc22]. Anticipations [GT23a]. Approach [Bis23, Tou22]. Archive [Huf21a, Huf21b, Huf22a, Huf22b, Sac23b, TGG21, Tou23]. Area [Cur21b, Cur21a]. Arteries [Bis21d]. Attracting [Bis21c].

Based [Tou22, BKMP22, BKR22]. Basel [HH22]. Binary [CW23]. Blood [Bis21d]. boats [Bis22d]. Bodies [Bis21c, Cur21a]. body [Bis21b].

Calculating [Hea22]. Canada [D’A21].

Caractère [BF21]. Change [GT22a, Bra22]. circa [Ayc22]. Class [Bis22b, Bis22c]. Collecting [TGG21]. Collection [BKMP22, BKR22]. Computandi [Hea22]. concerning [Ayc22]. Cone [Cur21b]. Cones [Cur21a]. Conical [Cur21a]. considerationes [Ayc22]. Contemporaries [GT21]. Contributions [TG21]. Conversations [GT22a]. Correspondence [Kly22].

David [D’A21]. Determining [Bis21d]. Difference [Ayc23d]. Digit [CW23]. Diophante [GS21]. Diophantine [Ehl23]. Duplication [Ayc23a].

E326 [Huf23]. E421 [Ayc21b]. Each [Bis21c]. Ecliptic [Hea23]. Eclipticam [Hea23]. Ed [Bra21]. Eiusque [Hea23]. Elastica [Bis23]. English [BF21, Gal21, Hea22, Hea23, Sac23b]. Equation [Ayc23d, Hea22, Ehl23]. Euler [Ayc21a, Ayc21b, Ayc22, Ayc23a, Ayc23c, Ayc23b, Ayc23d, BKMP22, BKR22, Bis21a, Bis21b, Bis22a, Bis23, BF21, CW23, Cur21b, Dun21, Gal21, GT21, GT23a, GT23b, Kly22, Huf21a, Huf21b, Huf22a, Huf22b, Sac23b, Sac23a, TGG21, Tou21, TG21, Tou23]. Euler-based [BKMP22, BKR22]. Eulerian [Bra21]. Extraction [CW23].

Father [Bis21a]. First [Ayc23c, CW23]. Formula [Ayc23a, Ayc23c, CW23, Ayc21a]. Found [CW23]. Function [Ayc21a, Ayc23a]. further [HH22].

Gamma [Ayc21a, Ayc23a]. Gamma-Function [Ayc23a].

Haemodynamics [Bis21a]. Historical [HH22]. History [D’A21, GS21, TGG21, Tou22]. hypergeometricas [Ayc22].

II [BKR22]. Inclination [Hea23]. Inclinationis [Hea23].

Lagrange [GS21]. Lambertina [Gal21]. Languages [Sac23b]. Latin [Gal21]. Learning [BKMP22, BKR22]. Legendre [Ayc23b]. Leibniz [Cur21b]. Leonhard [Kly22]. Look [GT21]. Lunae [Hea23].

Marathoner [Bra21]. Master [BKMP22, BKR22]. Mathematics [BKMP22, BKR22, D’A21, Tou22]. Meridiei [Hea22]. Method [Hea22]. Methodus [Hea22]. Miracle [Dun21]. Moderne [BF21]. Moon [Sac23a, Hea23]. Motion [Bis21d, Bis21c, Hea23, Bis22d]. Motu [Hea23]. multiplication [Ayc21a]. Musical [BF21]. Musique [BF21]. Mutually [Bis21c].

Navigation [Bis22a]. Nodes [Hea23]. Nodorum [Hea23]. Noon [Hea22]. Notes [Gal21]. Numbers [Has23, SW22, Ehl23].

oars [Bis22d]. Observations [Huf23]. Oscillations [Bis22b, Bis22c]. Other [Bis21c, Cur21a, Sac23b].

Paper [Ayc21b, Ayc22]. Part [Tou22, BKMP22, BKR22]. Perfect [Has23]. perspective [HH22]. Polynomials [Ayc23b]. Primary [BKMP22, BKR22]. Principles [Bis21d]. Problem [Bis22a, HH22, SW22, Bis21b]. problèmes [GS21]. processes [HH22]. Product [SW22]. Products [SW22]. Projects [BKMP22, BKR22]. Proof [Ayc23c]. proofs [HH22]. propelled [Bis22d]. Publishing [Tou21].

quelques [GS21]. question [Ayc22].

rational [Ehl23]. Rectilinear [Bis21c]. Review [Bra22, D’A21, Tou22]. rivers [Bis22d].

Sandifer [Bra21]. Scalene [Cur21b, Cur21a]. Schedule [Kly22]. Scholarship [TG21]. Scope [GT23b]. Searching [SW22]. Serie [Gal21]. series [Ayc22]. Seven [Tou21]. Sharing [TG21]. Simple [Ayc23d]. Solution [Ayc23d, Ehl23, SW22]. Solved [Cur21b]. Source [BKMP22, BKR22, Tou22]. Source-Based [Tou22]. Spotlight [Huf21a, Huf21b, Huf22a, Huf22b, Sac23b, Tou23]. Square [SW22]. States [D’A21]. Stirling [Ayc23c]. stochastic [HH22]. Students [BKMP22, BKR22]. Sum [SW22]. Supplementary [Gal21]. Surface [Cur21b, Cur21a]. Suspicious [Sac23a].

Taken [SW22]. Their [SW22]. Theories [BF21]. Three [Bis21c, SW22, Bis21b]. three-body [Bis21b]. Time [SW22]. Today [BKMP22, BKR22]. Translated [Gal21]. Translation [Ayc21b, BF21, GS21, Huf23, Hea22, Hea23]. Translations [Sac23b]. Tuning [BF21]. Two [SW22].

United [D’A21]. using [Ehl23].

Variae [Ayc22]. Variation [GT22b, Hea23]. Variational [Bis22a, Bis23]. Variatione [Hea23]. Variations [Bra22]. Variety [GT22b]. Varignon [Cur21b]. Venus [Sac23a]. Véritable [BF21]. Vol [D’A21, Tou22].

War [Tou21]. Wartime [Tou21]. Which [SW22]. Wide [GT23b]. Work [GT23b]. Works [Sac23b, TGG21].

Year [Tou21].

Zitarelli [D’A21].

References

Aycock:2021:EMF

[Ayc21a] Alexander Aycock. Euler and the multiplication formula for the gamma function. Euleriana, 1(2):197–204, August 2021. CODEN ???? ISSN 2693-9908. URL https://

scholarlycommons.pacific.edu/
euleriana/vol1/iss2/8.

Aycock:2021:TEP

[Ayc21b] Alexander Aycock. Translation of Euler’s paper E421. Euleriana, 1(2):205–251, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/9.

Aycock:2022:AQC

[Ayc22] Alexander Aycock. Answer to a question concerning Euler’s paper “Variae considerationes circa series hypergeometricas”. Euleriana, 2(2):113–119, September 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss2/7.

Aycock:2023:EDF

[Ayc23a] Alexander Aycock. Euler and the duplication formula for the gamma-function. Euleriana, 3(1): 31–35, March 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss1/4.

Aycock:2023:ELP

[Ayc23b] Alexander Aycock. Euler and the Legendre polynomials. Euleriana, 3(2):??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/4.

Aycock:2023:EFP

[Ayc23c] Alexander Aycock. Euler’s first proof of Stirling’s formula. Euleriana, 3(2):??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/6.

Aycock:2023:ESS

[Ayc23d] Alexander Aycock. On Euler’s solution of the simple difference equation. Euleriana, 3(2):

??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/5.

Blaine:2021:ETM

[BF21] Larry G. Blaine and Susan Ferré.

Euler’s theories of musical tuning with an English translation of Du Véritable Caractère de la Musique Moderne. Euleriana, 1(1):112– 140, February 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss1/8.

Bistafa:2021:EFH

[Bis21a] Sylvio R. Bistafa. Euler, father of haemodynamics. Euleriana, 1(1):88–92, February 2021. CODEN ???? ISSN 2693-9908. URL https://

scholarlycommons.pacific.edu/
euleriana/vol1/iss1/6.

Bistafa:2021:ETB

[Bis21b] Sylvio R. Bistafa. Euler’s three-body problem. Euleriana, 1(2):181–187, August 2021. CODEN ???? ISSN 2693-9908. URL https://

scholarlycommons.pacific.edu/
euleriana/vol1/iss2/6.

Bistafa:2021:RMT

[Bis21c] Sylvio R. Bistafa. On the rectilinear motion of three bodies mutually attracting each other. Euleriana, 1(2):188–196, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/7.

Bistafa:2021:PDM

[Bis21d] Sylvio R. Bistafa. Principles for determining the motion of blood through arteries. Euleriana, 1(1):93–111, February 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss1/7.

Bistafa:2022:ENV

[Bis22a] Sylvio R. Bistafa. Euler’s navigation variational problem. Euleriana, 2(2):131–142, September 2022. CODEN ???? ISSN 2693-9908. URL https://

                                        scholarlycommons.pacific.edu/
                                        euleriana/vol2/iss2/9.

Bistafa:2022:NCOa

[Bis22b] Sylvio R. Bistafa. On a new class of oscillations. Euleriana, 2(1):6– 10, March 2022. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss1/4.

Bistafa:2022:NCOb

[Bis22c] Sylvio R. Bistafa. On a new class of oscillations. Euleriana, 2(1):11– 26, March 2022. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss1/3.

Bistafa:2022:MBP

[Bis22d] Sylvio R. Bistafa. On the motion of boats propelled by oars in rivers. Euleriana, 2(2):79–93, September 2022. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss2/4.

Bistafa:2023:EVA

[Bis23]  Sylvio R. Bistafa.  Euler’s

variational approach to the elastica. Euleriana, 3(2):??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/7.

Barnett:2022:LMM

[BKMP22] Janet Heine Barnett, Dominic Klyve, Kenneth M. Monks, and Adam E. Parker. Learning mathematics from the master: A collection of Euler-based primary source projects for today’s students, Part I. Euleriana, 2(1):40– 50, March 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss1/6.

Barnett:2022:LMMb

[BKR22] Janet H. Barnett, Dominic Klyve, and Dave Ruch. Learning mathematics from the master: A collection of Euler-based primary source projects for today’s students, Part II. Euleriana, 2(2):94–106, September 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss2/5.

Bradley:2021:ESE

[Bra21] Robert E. Bradley. Ed Sandifer: An Eulerian marathoner. Euleriana, 1(2):165–167, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/3.

Bradley:2022:RCV

[Bra22] Robert E. Bradley. Review of Change and Variations. Euleriana, 2(1):3–5, March 2022. CO-DEN ???? ISSN 2693-9908. URL

                                        https://scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss1/2.

Curtin:2021:SASb

[Cur21a] Daniel J. Curtin. On the surface area of scalene cones and other conical bodies. Euleriana, 1(1):42–61, February 2021. CODEN ???? ISSN 2693-9908. URL https://

                                        scholarlycommons.pacific.edu/
                                        euleriana/vol1/iss1/4.

Curtin:2021:SASa

[Cur21b] Daniel J. Curtin. The surface area of a scalene cone as solved by Varignon, Leibniz, and Euler. Euleriana, 1(1):10–41, February 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss1/3.

Craig-Wood:2023:EFF

[CW23] Nick Craig-Wood. Euler found the first binary digit extraction formula for π in 1779. Euleriana, 3(1):23–30, March 2023. CODEN ???? ISSN 2693-9908. URL https://

                                        scholarlycommons.pacific.edu/
                                        euleriana/vol3/iss1/3.

DAntonio:2021:RHM

[D’A21] Lawrence D’Antonio. Review: A History of Mathematics in the United States and Canada (Vol. 1), by David Zitarelli.

Euleriana, 1(2):159–164, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/2.

Dunham:2021:EM

[Dun21] William Dunham. Euler’s miracle. Euleriana, 1(2):172–180, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/5.

Ehlers:2023:SDE

[Ehl23] Georg Ehlers. Solution of the Diophantine equation (maa + nbb) = cd(mcc + ndd) using rational numbers. Euleriana, 3(2): ??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/3.

Gallagher:2021:ESL

[Gal21] Sam Gallagher. Euler’s De Serie Lambertina, translated from Latin to English with supplementary notes. Euleriana, 1(2):252– 272, August 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss2/10.

Goff:2021:HTL

[GS21] Christopher Goff and Michael Saclolo. A history and translation of Lagrange’s “Sur quelques problèmes de l’analyse de Diophante”. Euleriana, 1(1):62–87, February 2021. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol1/iss1/5.

Goff:2021:NLE

[GT21] Christopher Goff and Erik Tou. A new look at Euler and his contemporaries. Euleriana, 1(1):1– 3, February 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss1/1.

Goff:2022:CC

[GT22a] Christopher Goff and Erik Tou. Conversations on change. Euleriana, 2(1):1–2, March 2022. CO-DEN ???? ISSN 2693-9908. URL

                                        https://scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss1/1.

Goff:2022:VV

[GT22b] Christopher Goff and Erik R. Tou. Variety and variation. Euleriana, 2(2):60–61, September 2022. CODEN ???? ISSN 2693-9908. URL https://

                                        scholarlycommons.pacific.edu/
                                        euleriana/vol2/iss2/1.

Goff:2023:EA

[GT23a] Christopher Goff and Erik Tou. Euler’s anticipations. Euleriana, 3(1):1–2, March 2023. CODEN ???? ISSN 2693-9908. URL

                                        https://scholarlycommons.pacific.
                                        edu/euleriana/vol3/iss1/1.

Goff:2023:WSE

[GT23b] Christopher Goff and Erik R. Tou. The wide scope of Euler’s work. Euleriana, 3(2): ??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol3/iss2/1.

Hassler:2023:PN

[Has23] Uwe Hassler. Perfect numbers. Euleriana, 3(2):??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https://

scholarlycommons.pacific.edu/
euleriana/vol3/iss2/8.

Headley:2022:MCE

[Hea22] Patrick T. Headley. A method for calculating the equation of noon (an English translation of Methodus Computandi Aequationem Meridiei). Euleriana, 2(2):67–78, September 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss2/3.

Headley:2023:MNM

[Hea23] Patrick T. Headley. On the motion of the nodes of the Moon and the variation of its inclination to the ecliptic (an English translation of De Motu Nodorum Lunae Eiusque Inclinationis Ad Eclipticam Variatione). Euleriana, 3(2): ??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/2.

Hassler:2022:BPH

[HH22] Uwe Hassler and Mehdi Hosseinkouchack. Basel problem: Historical perspective and further proofs from stochastic processes. Euleriana, 2(2):120–130, September 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss2/8.

PhD:2021:EAS

[Huf21a] Cynthia J. Huffman. Euler Archive spotlight. Euleriana, 1(1):141–143, February 2021. CODEN ???? ISSN 2693-9908. URL https://

                                        scholarlycommons.pacific.edu/
                                        euleriana/vol1/iss1/9.

PhD:2021:SEA

[Huf21b] Cynthia J. Huffman. Spotlight on the Euler Archive. Euleriana, 1(2):168–171, August 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss2/4.

PhD:2022:EASa

[Huf22a] Cynthia Huffman. Euler Archive spotlight. Euleriana, 2(1):56– 59, March 2022. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss1/8.

PhD:2022:EASb

[Huf22b] Cynthia Huffman. Euler Archive spotlight. Euleriana, 2(2):107– 112, September 2022. CODEN ???? ISSN 2693-9908. URL

                                        https://scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss2/6.

PhD:2023:AOT

[Huf23] Cynthia Huffman. Analytical observations (translation of E326). Euleriana, 3(1):3–22, March 2023. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol3/iss1/2.

Klyve:2022:LEC

[Kly22]  Dominic Klyve.  Leonhard

Euler’s correspondence schedule. Euleriana, 2(1):51–55, March 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss1/7.

Saclolo:2023:EVS

[Sac23a] Michael P. Saclolo. Euler and Venus’ suspicious moon. Euleriana, 3(1):36–41, March 2023. CODEN ???? ISSN 2693-9908. URL https://

scholarlycommons.pacific.edu/
euleriana/vol3/iss1/5.

Saclolo:2023:EAS

[Sac23b] Michael P. Saclolo. Euler Archive spotlight: Translations of Euler’s works to languages other than English. Euleriana, 3(2): ??, ???? 2023. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss2/9.

Snavely:2022:SPS

[SW22] Mark R. Snavely and Philip Woodruff. The solution of a problem of searching for three numbers, of which the sum, product, and the sum of their products taken two at a time, are square numbers. Euleriana, 2(1): 27–39, March 2022. CODEN ???? ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol2/iss1/5.

Tou:2021:SCE

[TG21] Erik R. Tou and Christopher Goff. Sharing contributions to Euler scholarship. Euleriana, 1(2):157– 158, August 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss2/1.

Tou:2021:CWH

[TGG21] Erik R. Tou, Christopher Goff, and Michele Gibney. Collecting works: A history of the Euler Archive. Euleriana, 1(1):4– 9, February 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss1/2.

Tou:2021:EWP

[Tou21] Erik R. Tou. Euler in wartime: Publishing in the Seven Years’ War. Euleriana, 1(1):144–156, February 2021. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol1/iss1/10.

Tou:2022:RHM

[Tou22] Erik R. Tou. Review of The History of Mathematics: A Source-Based Approach (Vol. 2), Part I. Euleriana, 2(2):62–66, September 2022. CODEN ???? ISSN 2693-9908. URL https:

                                        //scholarlycommons.pacific.
                                        edu/euleriana/vol2/iss2/2.

Tou:2023:EAS

[Tou23] Erik R. Tou. Euler Archive spotlight. Euleriana, 3(1):42– 44, March 2023. CODEN ????

ISSN 2693-9908. URL https:

//scholarlycommons.pacific.
edu/euleriana/vol3/iss1/6.